narutouzamaki2:

Fifty-four is blank % of 60.

narutouzamaki2:

plz help

supie:

So what we are doing is 54 % of 60

supie:

So \[54/100(60)=\]\[(54/100)(60/1)=\]\[(27/50)(60/1)=\]\[(27)(60)/ (50)(1)=\]\[1620/50=?\]

supie:

So since you said it makes no sense in the DM's, what I did was change the precent into a fraction.

supie:

So what you have to do is simplify the fraction \[1620/50=\] then you will get a simplified fraction. You have to change that into a decimal (precent) and then you get your answer.

supie:

Do you get it now @narutouzamaki2?

supie:

The question has already been answered. Why did you bump it? You can close it now unless you don't understand.

mnjones12:

its 54

narutouzamaki2:

thx

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mnjones12 its 54 \(\color{#0cbb34}{\text{End of Quote}}\) That's a direct answer and no.

supie:

54 isn't 54 precent of 60

mnjones12:

then what you think it is SUPIE